if the straight line with equation 12x-8y+24=0 and 4x+ky+16=0 are parallel, the value of k is:
a, 2
b, 1
c, -2
d, -4
e, -8/3

Answer by mananth(16946)

12x-8y+24=0 and 4x+ky+16=0 are parallel, the value of k is:
The slopes of parallel lines are equal
12x-8y+24=0
slope = -a/b = -12/-8=3/2
and 4x+ky+16=0
slope = -4/k
-4/k = 3/2
so k= -8/3